ย Show that the Jones matrix representing a linear polarizer whose transmission axis is at an arbitrary angle ๐œƒ with respect to the horizontal is given by


ย ๐‘€(๐œƒ) =( ย ๐‘๐‘œ๐‘ 2๐œƒย ย ย ย ย ย ย ย ๐‘ ๐‘–๐‘›๐œƒ๐‘๐‘œ๐‘ ๐œƒย ย ย ย )ย ย ย ย ย ย ย ย ย ย 

ย ย ย ย  ย ๐‘ ๐‘–๐‘›๐œƒ๐‘๐‘œ๐‘ ๐œƒ ย ๐‘ ๐‘–๐‘›2๐œƒย 


1
Expert's answer
2021-04-26T17:09:12-0400

The polarizer acts on a plane wave with arbitrary polarization. The electric field of our plane wave may be written as


"E(z,t)=(E_x \hat x+E_y\hat y)e^{i(kz-\omega t)}"

Let the transmission axis of the polarizer be specified by the unit vector "\hat e_1" and the absorption axis of the polarizer be specified by "\hat e_2" (orthogonal to the transmission axis). The vector "\hat e_1" is oriented at an angle ฮธ from the x-axis. We need to write the electric field components in terms of the new basis specified by "\hat e_1" and "\hat e_2". By inspection of the geometry, the x-y unit vectors are connected to the new coordinate system via:


"\hat x=cos\theta \hat e_1- sin\theta \hat e_2 \\\hat y=sin\theta \hat e_1+ cos\theta \hat e_2"

So, the electric field will be


"E(z,t)=(E_1 \hat e_1+E_2\hat e_2)e^{i(kz-\omega t)}"

After traversing the polarizer, the field becomes


"E_{after}(z,t)=(E_1 \hat e_1+\xi E_2\hat e_2)e^{i(kz-\omega t)}"

We now have the field after the polarizer, but it would be nice to rewrite it in terms of the original xโ€“y basis.


"\hat e_1=cos\theta \hat x- sin\theta \hat y \\\hat e_2=sin\theta \hat x+ cos\theta \hat y"

Substitution of these relationships into together with the definitionsย for "E_1" and "E_2" yields

"E_{after} (z,t) = [(E_xcos\theta+E_ysin\theta)(cos\theta\hat x+sin\theta\hat y)+ \xi[(-E_xsin\theta+E_ycos\theta)(-sin\theta\hat x+cos\theta\hat y)e^{i(kz-\omega t)}"


If we represent the electric field as a two-dimensional column vector with its x component in the top and its y component in the bottom (like a Jones vector), then we can rewrite


"E_{after}(z,t)=\begin{bmatrix}
 cos^2\theta+\xi sin^2\theta & sin\theta cos\theta-\xi sin\theta cos\theta \\
 sin\theta cos\theta-\xi sin\theta cos\theta & sin^2\theta+\xi cos^2\theta
\end{bmatrix}\begin{bmatrix}
 E_x \\
 E_y \end{bmatrix}e^{i(kz-\omega t)}"


The matrix here is a properly normalized Jones matrix, even though we did not bother factoring out "E_{eff}" to make a properly normalized Jones vector. We can now write down the Jones matrix for a polarizer by inserting ฮพ = 0 into the matrix:


"M(\theta)=\begin{bmatrix}
 cos^2\theta & sin\theta cos\theta \\
 sin\theta cos\theta & sin^2\theta
\end{bmatrix}" (polarizer with transmission axis at angle ฮธ)